OpenStudy (anonymous):

Please help!! Write the sum using summation notation, assuming the suggested pattern continues. -9 - 4 + 1 + 6 + ... + 66

OpenStudy (anonymous):

so you start at -9 and roll to 66 at steps of 5

OpenStudy (anonymous):

Ok. So would the answer be: summation of the quantity negative nine plus five n from n equals zero to infinity

OpenStudy (anonymous):

\[\sum_{n=-9}^{\infty}\]

OpenStudy (anonymous):

all my answer choices have n=0, then infinity or 15 at the top, and on the side: -45n or (-9+5n). So I thought I would have: infinity at the top, n=0 at the bottom, and (-9+5n) on the side.

OpenStudy (anonymous):

\[\sum_{k=0}^{\infty} = -9+(k-1)5\]

OpenStudy (anonymous):

change the k to n

OpenStudy (anonymous):

wait... that would mean n=0 is -14

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} = -4+(n-1)5\]

OpenStudy (anonymous):

that's not an answer choice

OpenStudy (anonymous):

that way n=0 is -9

OpenStudy (anonymous):

what choice do you have?

OpenStudy (anonymous):

a. summation of negative forty five times n from n equals zero to fifteen b. summation of the quantity negative nine plus five n from n equals zero to fifteen c. summation of the quantity negative nine plus five n from n equals zero to infinity d. summation of negative forty five times n from n equals zero to infinity

OpenStudy (anonymous):

looks like c if you want -9 to infinity

OpenStudy (anonymous):

That's what I thought.

OpenStudy (anonymous):

it works too

OpenStudy (anonymous):

just tested it

OpenStudy (anonymous):

That wasn't correct. Could it be b?

OpenStudy (anonymous):

yea 0 to 15 gives to -9 to 66

OpenStudy (anonymous):

b* dayum i soo sleepy i saw 0 as b